Find the lateral area of a right cone with radius 5 in. and height 12 in. Give your answers in terms of pie
Answer by amarjeeth123(570)

Lateral surface area of a cone=πr√(r^2+h^2)
=π(5)√25+144
=π(5)√169
=π(5)(13)
=65π square inches.